# Rethinking the Integration of Prediction and Planning in Deep Learning-Based Automated Driving Systems: A Review

### v2

@@ 4.5 Benchmarks
Another major challenge in closed-loop evaluation is to realistically simulate the SVs. Simple simulators such as Highway-Env [262] rely on non-reactive agents that follow predefined trajectories [263]. In the non-reactive version of nuPlan, and Waymax, these trajectories are obtained from real-world recordings. However, if the EV behaves differently than the expert during recording, this can result in collisions caused by the SVs, as they do not react to the EV. Hence, this simulation is usually limited to short sequences, where the EV stays close to the expert’s trajectory [264]. Carla, nuPlan, and Waymax also provide a reactive driver model for the SV, such as the Intelligent Driver Model (IDM) [16, 17, 265, 266, 267, 268]. Another line of work focuses on learning-based methods to simulate realistic driving behavior [269, 270, 271, 272, 273, 274]. In Waymax [233], users can plug in their own model to control objects in the scene.
5 Integrating prediction and planning
Figure 5: Connection between the integration principles, ego-to-agent relationship, and cost function. Red arrows for the robot leader planning indicate that applying these cost function categories would be possible but lacks a thorough empirical evaluation. We highlight this as an interesting direction for future research.
After reviewing key aspects such as interfaces and architecture of planning and prediction components found in ADSs, we change perspective and discuss the properties of the IPPS. In particular, we will highlight the influence of design decisions in the IPPS on behavior in interactive scenarios. We will start by categorizing existing works based on which components the IPPS comprises (Sec. 5.1). Then, we will describe the implications for the interactive behavior which follow from design choices in modular integrated systems (Sec. 5.2). Subsequently, different concepts of safety and contingency in modular integrated systems are discussed in Sec. 5.3. Finally, Sec. 5.4 discusses possible combinations of these categories. An overview is provided in Fig. 5.
5.1 Integration Principles
In this work, we distinguish among three design principles for IPPS. An overview can be seen in Fig. 6.
Monolithic E2E planning describes methods where the IPPS consists of a single E2E planner, which maps the state inputs $\overline{\textbf{X}}_{\textrm{SV}}$, $X_{\textrm{EV}}$, and $I$ to the EV’s trajectory $Y_{\textrm{EV}}$. Thus, the SVs’ future behavior and the interactions among them, as well as between the SVs and the EV, are not modeled explicitly. However, the expert’s actions, which the E2E planner learns to imitate in training, are based on reasoning of how the future will unfold. By learning to imitate actions, the model could exhibit behavior that is implicitly based on such reasoning as well.
For instance, consider a scenario where the EV’s leading vehicle brakes. An expert or a human driver will base their driving decision on the likely future scenario that the safety margin will shrink and thus brake as well. Hence, an E2E planner that imitates the expert’s behavior makes a decision that is implicitly based on a predicted future.
Pioneered by Pomerlau [9], there is a vast body of literature on IPPS based on monolithic E2E planners [19, 116, 117, 118, 22, 62, 10, 63, 64, 23, 67, 122, 236, 24]. The main disadvantage is their black-box nature, which makes model introspection and safety verification difficult.
Interpretable E2E planning frames prediction as an auxiliary learning task [58, 61, 120, 125]. Explicit predictions are an additional model output trained jointly with the planning task. Usually, a shared backbone encodes features that are used for both tasks. Then, separate heads decode the respective output representations.
The additional objective provides a learning signal to the backbone. It acts as a regularization which can make the learning process more sample-efficient and can increase generalization capabilities [58, 124, 119] as it was previously demonstrated for other auxiliary tasks in prediction and planning like speed prediction [64], semantic mapping [275], and occupancy forecasting [66]. However, the additional learning objective results in a trade-off between the different tasks, usually resulting in a hyperparameter that balances the respective losses and has to be tuned empirically.
Compared to monolithic E2E planning, the explicit prediction adds interpretability and facilitates introspection. Nonetheless, both IPPS designs cannot provide safety guarantees. Hence, we attribute the increased driving performance to the regularizing effect of the additional objective.
Manual integration of a planner with a prediction module means that a separate subsystem is employed for each of the tasks. The interplay of both tasks is manually engineered based on domain knowledge. The most widely adopted integration is a sequential pipeline, where first the prediction module is inferred, and its output is fed to a subsequent planning module [70, 66, 69, 123]. Since this design is unable to reflect the impact of the EV’s plan on the SVs, an alternative is to swap the order of prediction and planning: First, candidate plans are generated, followed by a conditional prediction module. Then, the best plan for the EV is selected among the candidates [129, 134]. Alternative methods iterate between prediction and planning [276] or do both jointly [132]. In the following, we give a chronological overview of different models and highlight the integration design of prediction and planning.
(a) Monolithic E2E IPSS
(b) Interpretable E2E IPSS
(c) Manually Integrated IPSS
Figure 6: IPPS design principles that show how prediction and planning can be integrated. Monolithic E2E methods (a) use no explicit prediction, while interpretable E2E methods (b) use it as a regularization in training. Manually integrated IPPSs (c) combine a modular prediction with a modular planning method using hand-designed interfaces.
PRECOG [65] uses a probabilistic model for conditional multi-agent forecasting. It implements autoregressive trajectory decoding for EV and SVs jointly. At each time step, the state of all agents is updated and used as input for all other agents in the next update step.
PiP [129, 277] reverses the view of IPPS. Its core idea is that SVs behave differently depending on the EV’s actions. A trajectory generator provides a set of candidate plans independent of the SVs. Each candidate plan is then used to condition prediction. Finally, the optimal plan is selected using a cost function.
DSDNet [68] combines E2E and modular model design. Each of the sequential NN modules additionally has access to the high-dimensional features of the perception backbone, complying with the hypothesis [20] that scene context is important for robust planning. For each detected vehicle, a set of potential future trajectories is predicted. To select the plan for the EV from a sampled candidate set, a hand-crafted cost function is applied, which quantifies the collision probability of the EV candidate trajectories and the SVs’ predicted trajectories.
P3 [66], LookOut [69], and MP3 [70] use cost functions to link prediction and planning. P3 predicts an occupancy map, LookOut [69] employs predicted trajectories to evaluate different pre-defined EV candidate trajectories. MP3 [70] extends P3 into a mapless approach by predicting an online map. A ’dynamic state map’ comprising detections and prediction is used together with the intended route to evaluate potential trajectories.
SafetyNet [130] parallelly applies a monolithic E2E planner with implicit prediction and an explicit prediction module. Multiple collision checks between EV plan and SVs predictions are performed. If the initial plan is classified as unsafe, a hand-crafted fallback layer generates a lane-aligned trajectory. SafePathNet[131] improves upon SafetyNet [130] with a transformer module for joint prediction and planning. In a mixture-of-experts approach, multiple modes are predicted and ranked. The ranking is updated by checking for collisions with the most probable SVs’ predictions. Finally, the highest scoring EV mode is selected.
DIPP [7] jointly predicts the trajectories for all agents, including the EV, which is treated as an SV. The mode with the highest probability is selected for each vehicle. A differentiable nonlinear motion planner then performs iterative, local optimization under consideration of a kinematic model and a hand-crafted cost function.
Inspired by hierarchical game-theoretic frameworks [278, 279, 280, 281], GameFormer [132] models interaction between agents as a level-$k$ game [282, 283]. Thus, a transformer decoder models the interaction between all agents, including the EV, by iteratively updating the individual predictions based on the predicted behavior of all agents from the previous level. Hence, the model yields a joint prediction for all agents, including the EV.
UniAD [133] implements a modular system that is trained E2E based on the planning objective. Queries are used as the interface between the submodules so that the planning module can attend to agent-level features from previous network layers. Various works adapted or extended this approach: VAD [135] and Hydra-MDP [284] use the output of previous layers explicitly for collision and map compliance checks. PPAD [276] repeatedly iterates between prediction and planning based on the perception modules’ outputs and GenAD [285] adapts a generative decoder to regress SV and EV trajectories jointly.
TPP [134], DTPP [136], and MBAPPE [286] represent multiple future outcomes in scenario trees, where branches reflect combinations of SV predictions and EV plans. The best plan is obtained through a tree search.
Compared to other integration principles, manual integration requires more engineering effort but restricts the solution space in a meaningful way by incorporating prior knowledge. Often, manually integrated IPPS provide a high explainability and safer plans than E2E systems [287, 131].
The timelines in Fig. 3 show how early works relied mostly on implicit predictions in monolithic E2E designs. Starting with ChauffeurNet [58] and PRECOG [65], explicit predictions became more popular. More recently, manual integration of E2E differentiable modular components gained popularity [203, 133].
5.2 Ego-to-Agents Relationship
After reviewing the integration principles from a system architecture perspective, in the following, we will analyze the manual integration category from a different perspective, namely the intended ego-to-agents relationship. This is of special importance in highly interactive scenarios since the EV needs to base its driving decisions on the observed and expected behavior of surrounding agents. However, as [5] pointed out, the EV also needs to be aware that it can influence the behavior of others.
For instance, in the example in Fig. 7, the plan of the EV can also influence the behavior of the SV: If it approaches the intersection faster, the SV might decelerate and yield to the EV. To describe if and in which way the mutual interaction of the EV and surrounding traffic is taken into account in the planning step, [5] introduced the following four categories: robot leader planning, human leader planning, co-leader planning, and joint planning. In this context, human refers to the surrounding traffic, and robot refers to the self-driving vehicle. In the following, we briefly describe these categories and discuss existing work.
Robot leader planning. The EV plan is inferred based on the current state, and the prediction step for the whole environment is conditioned on it. Hence, the EV can effectively seek to make the SV react to its trajectory [88, 128, 127]. This can result in aggressive driving behavior. For instance, in the example in Fig. 7, the EV would reason that when it follows the fast-progressing plan, the observed SV will yield to it to prevent a collision.
Human leader planning is the opposite of robot leader planning. The plan is based on the forecasted behavior of SVs [61, 62, 70, 66]. Hence, the EV shows a reactive behavior. It does not model the influence of the EV’s plan on the SVs, which can lead to underconfident behavior [5]. This is the case for IPPS that predict and plan sequentially. In the unprotected right turn example (cf. Fig. 7), the EV tries to find a plan that is suitable given the two predicted SV behaviors without being aware that it can influence them. Consequently, it will favor the slower plan.
Joint planning describes systems that are aware that vehicles (including ego) interact with each other. The EV’s plan is obtained by global optimization over all agents [288]. Hence, the IPPS system deterministically approximates a joint objective [289] under the assumption that an optimal outcome exists. Hence, it erroneously assumes that the behavior model used for SVs is exactly known and that every traffic participant optimizes this same global objective. For example, if the EV in Fig. 7 squeezes in before the approaching SV, this may be optimal with respect to a reasonable global objective. Nonetheless, there is no guarantee that the SV will behave accordingly. Consequently, [5] demonstrates how this can lead to fatal errors.
Co-leader planning is the most complex paradigm. It models the influence of SVs’ potential future behaviors on the EV plan as well as their reaction to potential ego trajectories [290, 291, 5]. In contrast to the joint planning category, the surrounding agents’ behavior is not assumed to be deterministic. The EV has to consider this uncertainty in the planning step, e.g., by being prepared for multiple future outcomes and reacting accordingly. Additionally, co-leader planning expects the SV to make plans for different potential behaviors of the EV. The EV can exploit this by behaving in a way that makes the SVs reveal their intention, which in turn lets the EV make more confident plans. For instance, in the example in Fig. 7, the EV can try to slowly approach the intersection to provoke a reaction of the SV. This reaction can either indicate that the SV is going to make a turn, e.g., with a turn signal or by slowing down, or the opposite, i.e., accelerating, to indicate that the SV will continue straight. The former will enable the EV to confidently decide in favor of the faster-progressing plan. CfO [5] originally introduced this paradigm, leveraging normalizing flows to represent multi-agent behavior as a learned coupling of independent variables. However, the intricate nature of this architecture may pose obstacles to wider adoption. Furthermore, their benchmark is simplistic, encompassing only a few scenarios with hard-coded agent behaviors. TPP [134] presents a readily comprehensible tree-structured framework that implements co-leader planning. It involves sampling kinematically feasible trajectories for the EV and then conditioning prediction on each of them. An autoregressive formulation for prediction ensures causality: It can be seen as a way to roll out the scene with SV predictions and the EV plan such that both can only access the observation history up until the respective rollout step. Finally, the best EV trajectory proposal is selected with a dynamic programming algorithm. DTPP [136] extends this with a transformer-based conditional prediction module, which allows the forecast of all agents conditioned on multiple EV plans at once. It also replaced the dynamic program with a learnable cost function and applied node pruning to reduce computational burden. Still, the EV reaction to the prediction is taken over entirely by the cost function because the sampled ego-proposals are non-reactive with respect to the SV predictions.
Fig. 3 shows that no clear trend toward one of the paradigms can be observed. While Co-leader planning is theoretically the most capable paradigm, it has not yet been shown empirically if one of these four theoretical concepts is superior to the others and how it can be realized in the design of system architectures. This may also be related to a lack of comprehensive empirical benchmarking focused on interactive behavior.
5.3 Safety and Contingency
Figure 7: Demonstration of contingency on an unprotected right turn. Let’s assume two expected future behaviors of the observed SV, namely $Y^{(1)}_{\textrm{SV}}$ and $Y^{(2)}_{\textrm{SV}}$ with probabilities of $P(Y^{(1)}_{\textrm{SV}})=0.2$ and $P(Y^{(2)}_{\textrm{SV}})=0.8$. Two candidate plans for the EV are considered $Y^{(1)}_{\textrm{EV}}$ (purple) and $Y^{(2)}_{\textrm{EV}}$ (orange). While the purple plan is only safe if the unlikely prediction $Y^{(1)}_{\textrm{SV}}$ occurs, the orange plan is safe under all predictions. Thus, it is preferred under worst-case assumptions. When planning with marginalized predictions, the cost function has to balance between the high progress of the purple plan and the potential safety risk if the unlikely prediction $Y^{(1)}_{\textrm{SV}}$ occurs. A contingency planner could exploit that both candidate plans have a common short-term action. Hence, it could opt for the purple plan as it is aware that it could still re-plan and fall back to the orange one.
The ego-to-agents relationship categories discussed above highlighted that considering multiple potential future scenarios is essential to safety and contingency planning. In the following, we will discuss how this can be incorporated into the cost function component $h$ that we defined as part of the planning function $f=h(g(X_{\textrm{EV}},\overline{\textbf{X}}_{\textrm{SV}},I)$ (cf. Sec. 4.4). We will analyze the implications of different design choices on safety and contingency. To this end, we introduce the following notation: We denote the number of relevant future scenarios with $N_{s}$. These scenarios can be reflected by a multinomial distribution $P\left(\overline{\textbf{Y}}^{(i)}_{\textrm{SV}}\right),i=1,..,N_{s}$ describing the behavior of SVs in each scenario. Based on how this distribution is reflected in the design of the IPPS, we form the following three groups of existing methods regarding safety and contingency: Planning with marginalized predictions, worst-case planning, and contingency planning. They are briefly outlined in Fig. 7.
Planning with marginalized predictions describes IPPS that do not explicitly distinguish multiple future scenarios $N_{s}$. This means that the prediction $\overline{\textbf{Y}}_{\textrm{SV}}$ is marginalized over the future outcomes $\overline{\textbf{Y}}^{(j)}_{\textrm{SV}}$. This can be done explicitly, i.e., by integrating over the scenario probabilities $P\left(\overline{\textbf{Y}}^{(j)}_{\textrm{SV}}\right)$
$\overline{\textbf{Y}}_{\textrm{SV}}=\sum_{j}^{N_{s}}\overline{\textbf{Y}}^{(j)}_{\textrm{SV}}\cdot P\left(\overline{\textbf{Y}}^{(j)}_{\textrm{SV}}\right)$ (9)
or implicitly, i.e., by representing all possible scenarios in a joint representation instead of differentiating scenarios. For instance, IPPS that incorporate the prediction in their cost function $h$ and do not consider separate future outcomes when evaluating a proposal fall into this category [61, 70, 66]. Hence, their selection function takes the form
$h=\operatorname*{argmin}_{i}c(Y^{(i)}_{\textrm{EV}},\overline{\textbf{Y}}_{\textrm{SV}}).$ (10)
Similarly, this is the case for all systems in the monolithic E2E category (cf. Sec. 5.1[201, 20, 24, 9, 19], as they do not explicitly consider multiple future scenarios and hence – if at all – learn marginalized predictions. Broadly speaking, this means that the IPPS assumes that all $N_{s}$ scenarios will occur to a certain degree at the same time. As can be seen in the example of Fig. 7, the cost function needs to trade off unlikely but dangerous scenarios (e.g. collision) against very likely low-cost scenarios. This is safety-critical since planners need to be cautious, especially w.r.t. unlikely but dangerous events.
Worst-case planning refers to IPPS that are aware that multiple future outcomes $\overline{\textbf{Y}}^{(j)}_{\textrm{SV}}$ exist. In this category, no probabilities $P(\overline{\textbf{Y}}^{(j)}_{\textrm{SV}})$ are considered. Instead, each proposal is evaluated based on the worst-case scenario. The motivation behind this assumption is that the final output trajectory must be safe in each possible scenario. Consequently, the selection function $h$ can be formulated as
This paradigm strongly focuses on collision avoidance and can result in overly conservative behavior, as shown in Fig. 7. While this assumption is not a popular choice in DL-based prediction and planning, it is broadly applied for rule-based safety layers such as RSS [292] and can be seen as an intermediate step between planning with marginalized predictions and contingency planning.
Contingency planning prepares for the unknown future development of the scene by taking into account different future scenarios $\overline{\textbf{Y}}^{(j)}_{\textrm{SV}}$ and their probabilities $P(\overline{\textbf{Y}}^{(j)}_{\textrm{SV}})$. The resulting plan hedges against worst-case risks while enabling progress in expectation. Both the cost function-based and the hybrid planning paradigm can have these properties, and the selection function $h$ takes the general form
$h=f_{c}(Y^{(i)}_{\textrm{EV}},\overline{\textbf{Y}}^{(j)}_{\textrm{SV}},P(\overline{\textbf{Y}}^{(j)}_{\textrm{SV}}))\,.$ (12)
where $f_{c}$ is a contingent selection function.
For instance, [69, 287, 293] find a plan that is safe for scenarios on the short horizon (independently of their probability, i.e., worst-case) and also optimize the expected cost on the long horizon (marginalization). Thus,
where $c_{s}$ and $c_{l}$ consider the short-term and the long-term part of the proposal $Y^{(i)}_{\textrm{EV}}$, respectively. In the example in Fig. 7, this planner allows the most progress while still ensuring safety. CfO [5] further differentiates between passive contingency where the EV is contingent upon the SVs [69, 287, 293, 294] and active contingency where the EV also expects the SVs to be contingent upon it [290, 291, 295].
5.4 Possible Combinations
In the previous sections, we have described three ways to categorize the integration of prediction and planning methods, namely (1) integration principle, (2) ego-to-agents relationships, and (3) safety and contingency. In the following, we discuss possible combinations of these three dimensions. Fig. 5 shows an overview. Our key insight is that the categories we characterized describe the differences between IPPS methods on different levels. While the integration principle focuses on a high-level system architecture, the ego-to-agent relationship is mostly based on interactive behavior. The safety and contingency consideration is founded on the cost function, i.e., a specific design choice for the proposal selection. In the following, we highlight associations between these categories.
In case of monolithic E2E [9, 20, 24, 19, 117, 64, 10, 121, 23, 67, 122, 62, 22, 118, 116, 201] and interpretable E2E [58, 61, 296], no part of the system architecture explicitly reflects an intended interactive behavior. Hence, they cannot be assigned to the ego-to-agents relationship categories, and we determine their ego-to-agents relationship as unknown.
For instance, imagine that a model is trained on demonstrations from an expert who follows the co-leader planning paradigm and employs contingency planning. Given that the model excels at imitating the expert, it could be assumed that the model behaves in the same way. However, there is no guarantee for this, and it could be equally possible that the model is unable to reason effectively about SVs, resulting in a behavior similar to the robot leader paradigm.
Similarly, as E2E models do not use hand-crafted cost functions with a pre-defined structure, the safety and contingency categories described in Sec. 5.3 do not apply to them.
In contrast, inferring the ego-to-agents relationship of modular systems is often possible. Thus, we can further divide manually integrated architectures into robot-leader, human-leader, joint planning, and co-leader planning.
There is a plethora of manual integration methods following the human leader paradigm (cf. Fig. 3). By following a traditional sequential scheme of prediction and planning, [70, 66, 69, 120, 68, 130] do not model the influence of the ego plan on SVs. This can be combined with any of the three safety and contingency categories. Works like DSDNet [68], P3 [66], MP3 [70], UniAD [133], and  [297] leverage marginalized predictions.
While applying a worst-case cost function is theoretically possible, most works that can distinguish multiple future scenarios aim to make contingency plans [69, 287, 293].
Similarly, we argue that the robot leader paradigm can be combined with all three cost functions. Consider a simple robot leader model that first identifies potential ego plans and then predicts the future behavior of SVs conditioned on each plan. The final selection of the EV’s plan can be based on a cost function belonging to each of the three categories. However, existing works use specialized cost functions that do not follow the structure we outlined in Sec. 5.3, e.g.,[129, 134, 128, 88]. We want to highlight that systematically combining robot-leader architectures with corresponding cost functions could be a promising direction for future research. Especially contingent or worst-case cost functions could mitigate the inherent problem associated with robot leader planning, namely relying on SVs to react in potentially unreasonable ways for the interests of the EV.
As described in Sec. 5.2, the active contingency resulting from the co-leader planning paradigm implies that the EV makes contingency plans while being aware that SVs are contingent upon it. Hence, it naturally employs a cost function that supports contingency plans.
Finally, the joint planning paradigm relies on optimizing a joint cost function for all traffic participants in the scene. Thus, none of the safety and contingency concepts are applicable, and we denote such a cost function as joint cost.
6 Outlook and Conclusion
Based on the overview we gave on deep learning-based prediction, planning, and their integration in automated driving systems, we identify two core challenges for future research: System design and comprehensive benchmarking. After discussing both, we draw a final conclusion.
System Design. Employing a traditional, strictly sequential system consisting of perception, prediction, planning, and control is still a popular choice. Our survey suggests that this approach is not able to meet the requirements set for driving systems. Alternative methods integrate prediction and planning in a way that allows conditioning prediction on potential ego plans and vice versa. Nonetheless, it remains unclear which exact integration architecture is most effective. Especially in the increasingly popular field of interpretable end-to-end systems, it is unclear how prediction and planning should be integrated.
Comprehensive Benchmarking. We discussed different aspects of the integration of prediction and planning. However, no comprehensive empirical benchmark reproduces and analyzes their strengths and weaknesses. Such an overview would help to better understand the effects of different ego-to-agents relationships and safety/contingency paradigms. This requires simulation in realistic and highly interactive scenarios with realistic driver models for surrounding vehicles and expressive interaction metrics. Moreover, benchmarks must cover corner case scenarios and test the generalization to distributional shifts.
In this work, we surveyed and analyzed the integration of prediction and planning methods in automated driving systems based on a comprehensive overview of the individual tasks and respective methods. We described, proposed, and analyzed categories to compare integrated prediction and planning works and highlighted implications on safety and behavior. Finally, we pointed out promising directions for future research based on the identified gaps.

### v3

@@ VI-B5 Traffic Agent Simulation
Another major challenge in closed-loop evaluation is to realistically simulate the SVs. In NAVSIM [288], Waymax, and the non-reactive version of nuPlan, traffic agents follow predefined trajectories obtained from real-world recordings. However, if the EV behaves differently than the expert during recording, this can result in collisions caused by the SVs, as they do not react to the EV. Hence, this simulation is usually limited to short sequences, where the EV stays close to the expert’s trajectory [289]. Carla, nuPlan, and Waymax provide a reactive driver model for the SV, such as the Intelligent Driver Model (IDM) [16, 290, 291, 292, 293]. In highway-env, traffic agents can be simulated with a driving model based on IDM+MOBIL [17], which is capable of performing lane changes. Another line of work focuses on learning-based methods to simulate realistic driving behavior [294, 295, 296, 297, 298, 299]. In Waymax [262], users can plug in their own model to control objects in the scene.
VI-B6 Discussion
Despite recent advancements in scenario generation and traffic simulation, there are no prominent large-scale benchmarks to evaluate the interactive behavior of IPPSs. Real-world datasets commonly comprise a limited number of lane changes or unprotected turns, and synthetic datasets with hand-crafted scenarios don’t scale to the needs of comprehensive benchmarking. Moreover, leading benchmarks do not report metrics highlighting the strengths and weaknesses of different IPPS architectures concerning their interactive behavior. We emphasize this as an important avenue for future research and hope our work inspires researchers to tackle this problem.
2016201720182019202020212022202320241988ALVINN [9]Deepdriving [190]PAC [220]End-to-End-CNN [25]FCN-LSTM [206]HRI [221]CIL $\uparrow$ [175]Camera360 [179]Seg2WP $\downarrow$ [21]CAL [181]DeepImitative $\uparrow$ [63]DOCP [10]ChauffeurNet $\downarrow$ [56]DILP $\uparrow$ [64]CILRS [65]PRECOG $\downarrow$ [66]NMP [62]UD-CIL $\uparrow$ [22]JLTP $\downarrow$ [191]MFP $\downarrow$ [90]LBC [68]PiP [217]LSD [180]DSDNet $\uparrow$ [69]P3 $\downarrow$ [67]LookOut $\uparrow$ [70]MP3 $\downarrow$ [71]Transfuser [26]NEAT $\uparrow$ [176]UrbanDriver [23]SafetyNet $\downarrow$ [189]LAV [171]DIPP [7]PlanT $\uparrow$ [187]SafePathNet [188]DiffStack [145]TPP $\uparrow$ [185]GameFormer $\uparrow$ [35]UniAD $\downarrow$ [37]VAD $\downarrow$ [36]PDM-Open $\uparrow$ [183]DTPP [38]GenAD $\uparrow$ [235]VADv2 $\downarrow$ [147]System ArchitectureModularMonolithic E2EModular E2EIPPSImplicitHumanRobotJointCo-LeaderPlanning ParadigmRegressionCost FunctionHybridIntegrated Prediction and Planning Systems
Figure 7: Important contributions to DL-based planning. We highlight how prediction and planning are integrated. The nodes and stem of each entry encode different aspects, which are explained below the timeline. To avoid overlaps works marked with $\leftarrow$ and $\rightarrow$ were slightly moved into the past and future, respectively.
VI-C Challenges and Trends
In the following section, we discuss existing trends and future challenges in the field of IPPSs. Based on our overview of deep learning-based IPPSs and their components, we identify three challenges for future research: system design, comprehensive benchmarking, and vehicle-to-X communication. After discussing them, we draw a final conclusion.
VI-C1 Trends
A chronological overview and categorization of influential works in the field is given in Fig. 7. Clustering research trends based on this overview yields three main phases of IPPSs. Until 2019 mainly monolithic systems with implicit prediction and output regression were used. Following works since 2020 parted with regression and instead incorporated prior knowledge in interpretable cost functions. Most of them integrated unconditioned predictions into the cost function and thus followed a human leader approach, which can lead to overly cautious behavior. Many works have already modularized the monolithic system and are E2E-trainable. Notably, almost all current state-of-the-art systems are end-to-end differentiable. In a third phase, beginning at the end of 2022, the integration of prediction and planning is getting more complex. In addition to sequential systems with human leader integration, undirected systems with joint IPPSs and bidirectional systems with co-leader IPPSs are implemented more frequently. Hybrid planning is also used more frequently than before. After cost functions were initially used to evaluate fixed candidate plans, newer methods often regress the candidates based on the scene and then apply the cost function.
VI-C2 System Design
Our survey suggests that strictly sequential architectures cannot meet the requirements set for driving systems. Nonetheless, it remains unclear which integration architecture is most effective and if the theoretical advantages of bidirectional IPPSs can be realized in practice. Especially the interactive behavior has seen little attention so far. We believe that E2E differentiable bidirectional IPPSs offer great potential for future study.
VI-C3 Benchmarking in Interactive Scenarios
We discussed different aspects of various IPPS architectures and design paradigms. However, no comprehensive empirical benchmark reproduces and analyzes their strengths and weaknesses. Such an overview would help to better understand the effects of different system architectures. This requires simulation in realistic and highly interactive scenarios with realistic driver models for surrounding vehicles and expressive interaction metrics. In this regard, data-driven simulation is a fundamental direction for future research. While powerful generative models have already been applied to scenario generation, simulating and evaluating the interactive behavior requires appropriate reactive policies to simulate traffic agents. Moreover, benchmarks must cover corner case scenarios and test the generalization to distributional shifts.
VI-C4 Vehicle-to-X Communication
The systems described in this work leverage predictions to represent the anticipated intentions of other traffic participants. With vehicle-to-vehicle or vehicle-to-X communication, these intentions don’t need to be predicted but can be transmitted to surrounding agents. Thus, joint optimization methods (see Sec. IV-B) that can optimize several vehicles’ goals can improve traffic efficiency and safety. However, such systems also need to hedge against erroneous transmissions or even malicious attacks. Moreover, they still need to be able to interact with non-communicating human-driven vehicles. Consequently, extending IPPS architectures to incorporate this additional information is a promising direction for future research.
VI-D Conclusion
In this work, we surveyed and analyzed the integration of prediction and planning methods in automated driving systems based on a comprehensive overview of the individual tasks and respective methods. We described, proposed, and analyzed categories to compare integrated prediction and planning works and highlighted implications on safety and behavior. Based on this, we discussed the shortcomings of undirected and sequential IPPSs and highlighted the need for IPPSs that model bidirectional interaction. We hope that this way of approaching the architectures of IPPSs opens a new perspective to understanding their implications and capabilities with respect to interaction with traffic. Finally, we pointed out promising directions for future research based on the identified gaps.